# A Review of Graph-Powered Data Quality Applications for IoT Monitoring Sensor Networks

### v1

@@ 3.1 Graph Signal Processing
The GSP emerged to adapt the signal processing foundations to signals with irregular support such as graphs [2, 1, 10]. A graph signal is defined as the map $x{:\;}\mathcal{V}\rightarrow\mathbb{R}$, then the $i$-th sensor measurement at a given time instant can be expressed as $x_{i}(t){\in}\mathbb{R}$ and the collection of measurements of the network at a given instant as $\mathbf{x}(t){\in}\mathbb{R}^{N}$ and a collection of measurements for a period of time as $\mathbf{X}{=}[\mathbf{x}(t_{1}),\dots,\mathbf{x}(t_{T})]{\in}\mathbb{R}^{N\times T}$. One of the key components of this framework is the graph shift matrix $\mathbf{S}$ which describes the graph topology, and allows us to define the notion of graph shift, the graph discrete Fourier transform (GDFT) as well as the notion of signal smoothness among others. From now on, we simplify the notation of a graph signal at a given instant as $\mathbf{x}(t)\equiv\mathbf{x}$.
@@ 3.1.1 Shift Matrix
To be able to define graph filters, we need to adapt the notion of classical signal frequency spectrum to the graph domain. For this purpose, the GDFT is used to provide the frequency spectrum of a graph signal via the graph shift matrix eigendecomposition [53]. For undirected graphs, the $\mathbf{S}$ is a symmetric matrix and its eigendecomposition is given as:
@@
where the eigenvector matrix $\mathbf{U}{=}[\mathbf{u}_{1},\dots,\mathbf{u}_{N}]{\in}\mathbb{R}^{N\times N}$ defines the components of the Fourier modes and $,\boldsymbol{\Lambda}{=}diag([\lambda_{1},$$\dots,\lambda_{N}]){\in}\mathbb{R}^{N\times N}$ includes the frequencies defined by the eigenvalues. Slow-varying components are associated with small eigenvalues while fast-changing components are associated with large eigenvalues. In this way, the GDFT can be defined as [12]:
@@
    Graph based on prior information: the graph is constructed using some prior information. An example can be the use of the geographical distance, where a weighting function based on the nodes’ geographical distance is used [2]. For instance, the Gaussian radial basis function is commonly used to assign weights, $W_{ij}{=}e^{\frac{-d_{ij}}{2\sigma}}$, where $d_{ij}{\in}\mathbb{R}$ is the geographical distance between nodes $i$ and $j$. Then, a threshold $TH{\in}\mathbb{R}$ can be used to force a certain graph sparsity. Another commonly used approach to restricting the sparsity of the resulting graph is building a K-nearest approach, i.e., only keeping the K most important edges per node. Some other examples may include the construction of the adjacency matrix according to the connectivity between sensors, i.e., $A_{ij}{=}1$ iff sensors $i$ and $j$ have communication.
@@ 3.1.2 Signal Smoothness
This metric has been widely used in problems such as graph learning, outlier detection, and missing value imputation [59, 55]. In the case $\mathbf{S}{=}\mathbf{L}$ then $S(\mathbf{L},\mathbf{x})=\sum_{i,j{\in}\mathcal{V}}W_{ij}(x_{i}-x_{j})^{2}$ meaning that the smaller the Laplacian quadratic form the more smooth the signal is. Nevertheless, we must place special care when using this criterion since its minimum is achieved for a total disconnected graph $\mathcal{E}{=}\varnothing$ or a null graph signal $\mathbf{x}{=}\mathbf{0}$. Analogously to classic signal processing, the graph TV can be defined through the definition of the graph shift operator $\mathbf{S}$ [60]:
@@
where $\epsilon{\in}\mathbb{R}$ and $\beta{\in}\mathbb{R}$ are hyperparameters. In fact, the minimization of a graph signal smoothness can be translated in terms of the GDFT coefficients as:
$\min_{\mathbf{x}}\;\;S_{\beta,\epsilon}(\mathbf{L},\mathbf{x})_{\beta,\epsilon}=\mathbf{x}^{\mathsf{T}}(\mathbf{L}+\epsilon\mathbf{I})^{\beta}\mathbf{x}=\sum_{i=1}^{N}(\lambda_{i}+\epsilon)^{\beta}\hat{\mathbf{x}}_{i}^{2}$ (9)
Thus, as the equation above denotes, the minimization of a graph signal smoothness can be translated to the penalization of the GDFT components $\hat{\mathbf{x}}_{i}$ according to the frequencies $\lambda_{i}$, being the largest frequencies the most penalized ones.
@@ 3.1.3 Graph Filters
Now that we have explained the use of the graph shift matrix as well as the utility of its eigendecomposition to formulate the GDFT, we can formulate the graph filters. We assume that the eigenvectors $\mathbf{u}_{i}$ are ordered in ascending order according to the eigenvalues, $\lambda_{1}{\leq}\dots\leq\lambda_{N}$. Given the definition of the GDFT we have that [12]:
$\mathbf{x}=\mathbf{U}\hat{\mathbf{x}}=\mathbf{U}\mathbf{U}^{-1}\mathbf{x}$ (10)
A non-parametric filter can be defined as a function of the eigenvalues or a function of the shift matrix [60]:
@@
To approximate the computation of the polynomials of $\mathbf{L}$ and $\boldsymbol{\Lambda}$ for a specific transfer function, different bases can be used, e.g., the Chebyshev polynomials, the Gegenbauer polynomials, and the monomial or Bernstein basis among others [67, 68]. These approximations provide a recursive solution to approximating the desirable transfer functions providing computational benefits. Most graph filters allow for a distributed implementation given the use of graph shifts, which can be seen as a result of a diffusion process where only connected nodes may share information [36]. This feature is particularly interesting in the field of sensor networks where edge and fog computing approaches can be employed to implement a specific data-driven model.
@@
$\tilde{\hat{\mathbf{x}}}=g(\boldsymbol{\Lambda})\mathbf{U}^{-1}\mathbf{x}=\begin{bmatrix}g(\lambda_{1})&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&g(\lambda_{N})\end{bmatrix}\mathbf{U}^{-1}\mathbf{x}$ (15)
@@ 3.2 Machine Learning over Graphs
The use of machine learning on data residing in a manifold represented by a graph has been in use for years [25, 33, 7]. In the field of graph-based semi-supervised learning, data is assumed to reside on a manifold, with data points being related to each other, so that models can input such structured information to assist machine learning models. Other examples of machine learning tasks include the inference of node embeddings from graph data [72]. In this context, instead of working directly with graph-structured data, node embeddings are obtained and classic machine learning models are fed with these node representations.
In this section, we delve into the main uses of machine learning on graphs in sensor networks, which are the development of graph-regularized models and the interpretability they provide. Although many tasks and techniques could be included in graph machine learning, e.g., obtaining metrics, graph representation learning, node embeddings, or community detection, in this paper we focus on its use mainly through regularization and kernels in predictive models given its potential applicability to data quality in sensor networks.
3.2.1 Graph-Regularized Machine Learning
In many cases, it is assumed that the data resides in a non-Euclidean space, such as a smooth manifold or a Riemannian manifold [7, 73]. Hence, a graph can be used to describe the similarities of the data in such a manifold. Graph machine learning models aim to take advantage of a graph that describes the relationship between the data. This concept has been widely used in graph semi-supervised learning, regularization, and kernels over graphs. Thus, a manifold $\mathcal{M}$ is approximated via a graph $\mathcal{G}$, Figure 8.
The Laplace-Beltrami operator of functionals on Riemannian manifolds is widely used and the graph Laplacian operator is its discrete analog when the number of data points tends to infinity [74, 75]. Then, the smoothness of a function over a graph is represented by the quadratic form [47, 7]:
$\mathbf{x}^{\mathsf{T}}\mathbf{L}\mathbf{x}=\sum_{i,j\in\mathcal{V}}W_{ij}(x_{i}-x_{j})^{2}$ (16)
@@ 3.2.1 Graph-Regularized Machine Learning
Now, we can illustrate the use of graph-based regularization in a regression task, independently of whether it is supervised or semi-supervised, as finding the map $f{:\;}\mathbb{R}^{M}\rightarrow\mathbb{R}^{N}$ where $M{\geq}1$:
@@
This is an example of a transductive method, where for a given graph signal with some observed values, the signal is reconstructed taking into account the resulting signal smoothness. This formulation has been widely used for the regression of graph signals in scenarios such as image processing and sensor networks [7, 38, 77]. Similarly, in the context of supervised machine learning and semi-supervised machine learning, the norm in the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$ associated to a graph $\mathcal{G}$ has been used in Laplacian regularized least squares and in Laplacian support vector machines [78], $R_{\mathcal{G}}(f){\equiv}\|f\|^{\mathcal{G}}_{\mathcal{H}}$.
More remarkably, the graph counterpart of RKHS has been used in graph machine learning [79, 80, 38, 81]. Graph kernels play an important role in machine learning for comparing graphs in tasks such as graph classification and sub-graph matching in fields such as biology [82, 83]. In sensor networks, however, we are interested in graph kernels that compare the relationships between the nodes in a graph, $k{:\;}\mathcal{V}\times\mathcal{V}\rightarrow\mathbb{R}$. In this sense, we can define a graph kernel matrix $\mathbf{K}_{\mathcal{G}}$ that depicts the similarities between graph nodes as [79]:
$\mathbf{K}_{\mathcal{G}}$ $=r(\mathbf{L})=\mathbf{U}r^{\dagger}(\boldsymbol{\Lambda})\mathbf{U}^{-1}$ (19)
where $r{:\;}\mathbb{R}\rightarrow\mathbb{R}_{+}$ is a non-negative map and $\dagger$ denotes the pseudoinverse. Similarity functions such as graph diffusion or random-walk kernels have been used for graphs. This concept led to the development of graph kernel-based regression where the RKHS $\mathcal{H}$ is defined for graph signals and its norm is used for regularization [79]:
$R_{\mathcal{G}}(f)=\|f\|_{\mathcal{H}}=<f,f>_{\mathcal{H}}=\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{K}_{\mathcal{G}}\boldsymbol{\alpha}$ (20)
@@
3.2.2 Interpretable Machine Learning
The interpretability of machine learning and artificial intelligence models has been the focus of research in recent years given the increasing development and adoption of highly complex models such as neural networks (commonly known as black-box models). The interpretability of a machine learning model aims to reason why a prediction is made and to give further insights into the performance and functioning of a model [90, 91].
Interpretable machine learning promotes techniques that are interpretable, such as generalized linear regression, generalized additive models, and decision trees among others [92]. Among the most common approaches to interpretability are; i) model-specific interpretability, i.e., interpreting the inferred parameters, such as the coefficients of a linear regression, ii) sparsity-promoting models to force feature selection and enhance the interpretability of the parameters, and iii) model-agnostic interpretability, i.e., providing black-box models with tools to explain the results, such as neural network unrolling, feature visualization, latent embedding inspection, or results inspection.
Signal processing and machine learning over graphs have been shown to enhance the interpretability of models [9]. First, a graph provides an explicit structure that indicates the relationships between the sensors as well as the importance of the relationships. Thus, the graph shift matrix defines a node’s neighborhood $\mathcal{N}(i){=}\{j{:\;}j{\in}\mathcal{V}\wedge S_{ij}{\neq}0\}$, i.e., the set of influential nodes for a given node. In this line, in graph attention networks (GAT) the attention mechanism defines the importance of the nodes’ features (or hidden features) in a similar manner as the graph shift matrix but allows attention at different levels, e.g., different layers and multi-head attention mechanism. The graph learning field seeks to identify the network structure, i.e., relationships between nodes, by promoting smoothness and frequency-domain constraints, therefore, improving the interpretability of the resulting graph. Furthermore, GSP has been used to enhance the interpretability of classic neural networks and graph neural networks by imposing graph-based regularization in hidden layers, forcing the hidden features to adapt to a given graph [93]. In this line, graph filters and graph signal denoising techniques provide an interpretation in the frequency domain, enhancing the interpretability of the results [9]. In fact, GNNs have been interpreted as the result of signal denoising and filtering processes [94]. Finally, it is worth noting that kernels over graphs and the use of kernelized graph methods also provide an interpretable alternative since the kernel map can denote similarities between nodes and similarities between graphs. The latter concept has been employed in the development of interpretable GNN using graph kernels [95].
@@ 3.3 Graph Neural Networks
More precisely, we can broadly classify GNNs into [26]; i) graph convolutional neural networks (GCNs), ii) graph autoencoders (GAEs), and iii) graph recurrent neural networks. Spatiotemporal GNNs are also frequently included in this classification since they incorporate different building modules and may have tailored architectures. In this review, we will mention their main components without going into detail. More precisely, first, we will describe the building blocks of GNNs; i) propagation modules, ii) sampling, iii) pooling, and iv) readout function. Then, we will describe the main GNN architectures listed above, placing special emphasis on GCNs and their components given that these are widely used and can be used to implement other GNN architectures such as autoencoders. In the next sections, we explain the core elements of the different GNNs, and in section 4 we review well-known architectures used for specific sensor network applications.

### v2

@@ 3.1 Graph Signal Processing
The GSP emerged to adapt the signal processing foundations to signals with irregular support such as graphs [2, 1, 10]. A graph signal is defined as the map $x{:\;}\mathcal{V}\rightarrow\mathbb{R}$, then the $i$-th sensor measurement at a given time instant can be expressed as $x_{i}(t){\in}\mathbb{R}$ and the collection of measurements of the network at a given instant as $\mathbf{x}(t){\in}\mathbb{R}^{N}$ and a collection of measurements for a period of time as $\mathbf{X}{=}[\mathbf{x}(t_{1}),\dots,\mathbf{x}(t_{T})]{\in}\mathbb{R}^{N\times T}$. One of the key components of this framework is the graph shift matrix $\mathbf{S}$ which describes the graph topology, and allows us to define the notion of graph shift, the GDFT, as well as the notion of signal smoothness among others. From now on, we simplify the notation of a graph signal at a given instant as $\mathbf{x}(t)\equiv\mathbf{x}$. Figure 7 illustrates the use of the GSP framework to model the measurements of a sensor network, where the measurements are defined as signals on the graph and the graph shift matrix $\mathbf{S}$ describes the relationships between the sensors.
Figure 7: The GSP framework defines sensor network measurements as signals defined over graphs.
@@ 3.1.1 Shift Matrix
To be able to define graph filters, we need to adapt the notion of classical signal frequency spectrum to the graph domain. For this purpose, the GDFT is used to provide the frequency spectrum of a graph signal via the graph shift matrix eigendecomposition [59]. The eigendecomposition of matrix $\mathbf{S}$ is defined as:
@@
where the eigenvector matrix $\mathbf{U}{=}[\mathbf{u}_{1},\dots,\mathbf{u}_{N}]{\in}\mathbb{R}^{N\times N}$ defines the components of the Fourier modes and $,\boldsymbol{\Lambda}{=}diag([\lambda_{1},$$\dots,\lambda_{N}]){\in}\mathbb{R}^{N\times N}$ includes the frequencies defined by the eigenvalues. For undirected graphs, the matrix $\mathbf{S}$ is real symmetric so $\mathbf{U}^{-1}{=}\mathbf{U}^{\mathsf{T}}$. Slow-varying components are associated with small eigenvalues while fast-changing components are associated with large eigenvalues. In this way, the GDFT can be defined as [12]:
@@
    Graph based on prior information: the graph is constructed using some prior information. An example can be the use of the geographical distance, where a weighting function based on the nodes’ geographical distance is used [2]. For instance, the Gaussian radial basis function is commonly used to assign weights, $W_{ij}{=}e^{\frac{-d_{ij}}{2\sigma}}$, where $d_{ij}{\in}\mathbb{R}$ is the geographical distance between nodes $i$ and $j$. Then, a threshold $TH{\in}\mathbb{R}$ can be used to force a certain graph sparsity. Another commonly used approach to restricting the sparsity of the resulting graph is building a K-nearest neighbor approach, i.e., only keeping the K most important edges per node. Some other examples may include the construction of the adjacency matrix according to the connectivity between sensors, i.e., $A_{ij}{=}1$ iff sensors $i$ and $j$ have communication.
@@ 3.1.2 Signal Smoothness
This metric has been widely used in problems such as graph learning, outlier detection, and missing value imputation [65, 61]. In the case $\mathbf{S}{=}\mathbf{L}$ then $S(\mathbf{L},\mathbf{x})=\frac{1}{2}\sum_{i,j{\in}\mathcal{V}}W_{ij}(x_{i}-x_{j})^{2}=$ meaning that the smaller the Laplacian quadratic form the more smooth the signal is. Nevertheless, we must place special care when using this criterion since its minimum is achieved for a total disconnected graph $\mathcal{E}{=}\varnothing$ or a constant or null graph signal $\mathbf{x}{=}\mathbf{0}$. Analogously to classic signal processing, the graph TV can be defined through the definition of the graph shift operator $\mathbf{S}$ [66]:
@@
where $\epsilon{\in}\mathbb{R}$ and $\beta{\in}\mathbb{R}$ are hyperparameters. In fact, the minimization of a graph signal smoothness can be translated to the penalization of the GDFT components $\hat{\mathbf{x}}_{i}$ according to the frequencies $\lambda_{i}$, being the largest frequencies the most penalized ones [68].
@@ 3.1.3 Graph Filters
Now that we have explained the use of the graph shift matrix as well as the utility of its eigendecomposition to formulate the GDFT, we can formulate the graph filters. We assume that the eigenvectors $\mathbf{u}_{i}$ are ordered in ascending order according to the eigenvalues, $\lambda_{1}{\leq}\dots\leq\lambda_{N}$. Given the definition of the GDFT [12], a non-parametric filter can be defined as a function of the eigenvalues in frequency domain, or a function of the shift matrix in vertex domain [66]:
@@
To approximate the computation of the polynomials of $\mathbf{L}$ and $\boldsymbol{\Lambda}$ for a specific transfer function, different bases can be used, e.g., the Chebyshev polynomials, the Gegenbauer polynomials, and the monomial or Bernstein basis among others [73, 74, 75]. These approximations provide a recursive solution to approximating the desirable transfer functions providing computational benefits. Most graph filters allow for a distributed implementation given the use of graph shifts, which can be seen as a result of a diffusion process where only connected nodes may share information [40]. This feature is particularly interesting in the field of sensor networks where edge and fog computing approaches can be employed to implement a specific data-driven model.
@@
$\mathbf{x}\ast g=\mathbf{U}g(\boldsymbol{\Lambda})\mathbf{U}^{-1}\mathbf{x}=\mathbf{U}\begin{bmatrix}g(\lambda_{1})&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&g(\lambda_{N})\end{bmatrix}\mathbf{U}^{-1}\mathbf{x}$ (13)
@@ 3.2 Machine Learning over Graphs
The use of ML on data residing in a manifold represented by a graph has been in use for years [29, 37, 7]. In the field of graph-based semi-supervised learning, data is assumed to reside on a manifold, with data points being related to each other, so that models can input such structured information to assist ML models. Other examples of ML tasks include the inference of node embeddings from graph data [79]. In this context, instead of working directly with graph-structured data, node embeddings are obtained and classic ML models are fed with these node representations.
In this section, we delve into the main uses of ML on graphs in sensor networks, which are the development of graph-regularized models and the interpretability they provide. Although many tasks and techniques could be included in graph ML, e.g., obtaining metrics, graph representation learning, node embeddings, or community detection, in this paper we focus on its use mainly through regularization and kernels in predictive models given its potential applicability to data quality in sensor networks.
3.2.1 Graph-Regularized ML
In many cases, it is assumed that the data resides in a non-Euclidean space, such as a smooth manifold or a Riemannian manifold [7, 80]. Hence, a graph can be used to describe the similarities of the data in such a manifold. Graph ML models aim to take advantage of a graph that describes the relationship between the data. This concept has been widely used in graph semi-supervised learning, regularization, and kernels over graphs. Thus, a manifold $\mathcal{M}$ is approximated via a graph $\mathcal{G}$, Figure 9.
@@ 3.2.1 Graph-Regularized ML
The Laplace-Beltrami operator of functionals on Riemannian manifolds is widely used and the graph Laplacian operator can be seen as its discrete approximation when the number of data points tends to infinity [81, 82, 2]. Then, the smoothness of a function over a graph is represented by the quadratic form $\mathbf{x}^{\mathsf{T}}\mathbf{L}\mathbf{x}$, which is akin to the graph signal smoothness [52, 7]. Now, we can illustrate the use of graph-based regularization in a regression task, independently of whether it is supervised or semi-supervised, as finding the map $f{:\;}\mathbb{R}^{M}\rightarrow\mathbb{R}^{N}$ where $N{>}M{\geq}1$:
@@
This is an example of a transductive method, where for a given graph signal with some observed values, the signal is reconstructed taking into account the resulting signal smoothness. This formulation has been widely used for the regression of graph signals in scenarios such as image processing and sensor networks [7, 42, 84]. Similarly, in the context of supervised ML and semi-supervised ML, the norm in the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$ associated to a graph $\mathcal{G}$ has been used in Laplacian regularized least squares and in Laplacian support vector machines [85], $R_{\mathcal{G}}(f){\equiv}\|f\|^{\mathcal{G}}_{\mathcal{H}}$.
More remarkably, the graph counterpart of RKHS has been used in graph ML [86, 87, 42, 88]. Graph kernels play an important role in ML for comparing graphs in tasks such as graph classification and sub-graph matching in fields such as biology [89, 90]. In sensor networks, however, we are more interested in graph kernels that compare the relationships between nodes in a graph, $k{:\;}\mathcal{V}\times\mathcal{V}\rightarrow\mathbb{R}$. In this sense, we can define a graph kernel matrix $\mathbf{K}_{\mathcal{G}}$ that depicts the similarities between graph nodes as [86]:
$\mathbf{K}_{\mathcal{G}}$ $=r^{\dagger}(\mathbf{L})=\mathbf{U}r^{\dagger}(\boldsymbol{\Lambda})\mathbf{U}^{-1}$ (16)
where $r{:\;}\mathbb{R}\rightarrow\mathbb{R}_{+}$ is a non-negative map and $\dagger$ denotes the pseudoinverse. Similarity functions such as graph diffusion or random-walk kernels have been used for graphs. This concept led to the development of graph kernel-based regression where the RKHS, $\mathcal{H}$, is defined for graph signals and its norm is used for regularization [86]:
$R_{\mathcal{G}}(f)=\|f\|_{\mathcal{H}}^{2}=<f,f>_{\mathcal{H}}=\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{K}_{\mathcal{G}}\boldsymbol{\alpha}$ (17)
@@
3.2.2 Interpretable ML
The interpretability of ML and artificial intelligence models has been the focus of research in recent years given the increasing development and adoption of highly complex models such as neural networks (commonly known as black-box models). The interpretability of a ML model aims to reason why a prediction is made and to give further insights into the performance and functioning of a model [97, 98].
Interpretable ML promotes techniques that are interpretable, such as generalized linear regression, generalized additive models, and decision trees among others [99]. Among the most common approaches to interpretability are; i) model-specific interpretability, i.e., interpreting the inferred parameters, such as the coefficients of a linear regression, ii) sparsity-promoting models to force feature selection and enhance the interpretability of the parameters, and iii) model-agnostic interpretability, i.e., providing black-box models with tools to explain the results, such as neural network unrolling, feature visualization, latent embedding inspection, or results inspection.
Signal processing and ML over graphs have been shown to enhance the interpretability of models [9]. First, a graph provides an explicit structure that indicates the relationships between the sensors as well as the importance of the relationships. Thus, the graph shift matrix defines a node’s neighborhood $\mathcal{N}(i){=}\{j{:\;}j{\in}\mathcal{V}\wedge S_{ij}{\neq}0\}$, i.e., the set of influential nodes for a given node. In this line, in graph attention networks (GAT) the attention mechanism defines the importance of the nodes’ features (or hidden features) in a similar manner as the graph shift matrix but allows attention at different levels, e.g., different layers and multi-head attention mechanism. The graph learning field seeks to identify the network structure, i.e., relationships between nodes, by promoting smoothness and frequency-domain constraints, therefore, improving the interpretability of the resulting graph. Furthermore, GSP has been used to enhance the interpretability of classic neural networks and graph neural networks by imposing graph-based regularization in hidden layers, forcing the hidden features to adapt to a given graph [100]. In this line, graph filters and graph signal denoising techniques provide an interpretation in the frequency domain, enhancing the interpretability of the results [9]. In fact, GNNs have been interpreted as the result of signal denoising and filtering processes [101]. Finally, it is worth noting that kernels over graphs and the use of kernelized graph methods also provide an interpretable alternative since the kernel map can denote similarities between nodes and similarities between graphs. The latter concept has been employed in the development of interpretable GNN using graph kernels [102].
@@ 3.3 Graph Neural Networks
More precisely, we can broadly classify GNNs into [30]; i) graph convolutional neural networks (GCNs), ii) graph autoencoders (GAEs), and iii) graph recurrent neural networks. Spatiotemporal GNNs are also frequently included in this classification since they incorporate different building modules and may have tailored architectures. In this review, we will mention their main components without going into detail. More precisely, first, we will describe the building blocks of GNNs; i) propagation, ii) sampling, iii) pooling and readout functions. Then, we will describe the main GNN architectures listed above, placing special emphasis on GCNs and their components given that these are widely used and can be used to implement other GNN architectures such as autoencoders. In the next sections, we explain the core elements of the different GNNs, and in section 4 we review well-known architectures used for specific sensor network applications.